OpenStudy (unknownrandom):

Not sure how to do this? Use log5 2 = .4307 and log5 3=.6826 to approximate the value of each expression. log5 30

OpenStudy (unknownrandom):

@satellite73

OpenStudy (mathmale):

First: What is log 10? (Note: the base is 10) What is the value of \[\log_{5} 5=\log_{5}5^1? \] Supposing that we now have: \[\log_{5}2=0.4307 ,~\log_{5}3 =0.6826~and~\log_{5} 5=1.0000\] Now we're ready to find \[\log_{5}30. \] Ask yourself: How would you use the factors 2, 3 and 5 to obtain 30? Note: 2*3 = 6, right? Therefore, \[\log_{5} 3+\log_{5}2=\log_{5}6 \] think about this. If you know how to make use of the factors 2, 3 and 5 to obtain 30, then you can use a similar approach to find \[\log_{5}30 \]

OpenStudy (unknownrandom):

Thanks @mathmale! I think I understand now. I got 2.1133.

OpenStudy (mathmale):

You've done very well!! Try this on your calculator: 5^2.1133. What do you get? What does that tell you?

OpenStudy (unknownrandom):

I got 30.

OpenStudy (unknownrandom):

So they balance out. I am not sure what it tells me though?

OpenStudy (unknownrandom):

I know the answer is valid because they equal the same thing. Is that what you are asking?

OpenStudy (mathmale):

Exactly. You have correctly found the exponent of 5 that produces 30. Nice work!!

OpenStudy (unknownrandom):

I have a question on this one log5 (2/3)

OpenStudy (mathmale):

That's the log of a quotient. The rule of logs for that is \[\log \frac{ a }{ b }= \log a - \log b\], or the other way around: Can you apply tht rule to your "log5 (2/3)?"